What is the profitability index for an investment with the following cash flows given a 10 percent required return?

Year Cash Flow
0 $-33,000
1 $13,700
2 $16,000
3 $15,400

Note: Please explain how the present value factor is calculated.
Answer #1

Present value = 13,700 / (1 + 0.1)1 + 16,000 / (1 + 0.1)2 + 15,400 / (1 + 0.1)3

Present value = 13,700 / 1.1 + 16,000 / 1.21 + 15,400 / 1.331

Present value = 12,454.5455 + 13,223.1405 + 11,570.2479

Present value = $37,247.9339

Profitability index = Present value / initial investment

Profitability index = 37,247.9339 / 33,000

Profitability index = 1.13